OpenStudy (anonymous):

square root of 8 times square root of 96

OpenStudy (anonymous):

\[\sqrt{8}\times\sqrt{96}\]\[\sqrt{4\times2}\times\sqrt{16\times6}\]\[2\sqrt{2}\times 4\sqrt{6}\]\[8\sqrt{12}\]\[8\sqrt{4\times3}\]\[16\sqrt{3}\]